OpenStudy (anonymous):

WILL give MEDAL ! ! ! (g+4) - 3g = 1 + g a) g=1 b) g=4 c) no solution d) identity

OpenStudy (anonymous):

To solve for g, you need to combine all like terms

OpenStudy (anonymous):

So, (g+4)-3g=1+g g+4-3g=1+g -2g+4=1+g 3=3g g=1. So the answer is a

OpenStudy (asiyer):

g+4-3g=1+g 4-2g=1+g solving on both sides 4-1=3g 3=3g g=1

OpenStudy (anonymous):

Thanks so much for an explanation! Some people just say blah blah blah and just leave me hanging! Thanks again!!! <@_@>

OpenStudy (anonymous):

No problem!